A rectangular painting has a diagonal measurement of 13 inches and an area of 60 square inches. find the length and the width.
Answer by oscargut(2103)


Let x = width and y = length
xy = 60
x^2+y^2 = 13^2 = 169
Solving it:
Length = 12
Width = 5
You can ask me more at: mthman@gmail.com
Thanks